OpenStudy (anonymous):

Write the equation of the line that passes through (1, 3) and (4, 4) in standard form. 3x – 3y = –8 x – 3y = –9 x – 3y = –8 3x + y = 9

OpenStudy (ghazi):

y-4=(4-3)/(4-1)*(x-4)

OpenStudy (ghazi):

y-4= 1/3(x-4) 3y-12=x-4 x-3y=-8